OpenStudy (anonymous):

Suppose the perimeter of a square is 24. What is its area?

OpenStudy (ash2326):

Perimeter=sum of all sides=4a so 24=4a so side of square=6

OpenStudy (ash2326):

Area= a^2=6*6=36

OpenStudy (anonymous):

36 square units.